OpenStudy (anonymous):

Find the x intercepts of y=x^6-5x^5+8x^4-4x^3

OpenStudy (rmrjr22):

i assume you would have to factor out x to minimize the eqation

OpenStudy (rmrjr22):

y = x^3(x^3 -5x^2+8x-4)

OpenStudy (anonymous):

maybe factor out x^3 so it can have a constant

OpenStudy (anonymous):

better to use the rational root theorem after?

OpenStudy (rmrjr22):

you can still break it down

OpenStudy (rmrjr22):

x^3(x-2)^2(x-1) and set each portion to 0 and solve

OpenStudy (anonymous):

what happens with the rest of the terms outside the parenthesis?

OpenStudy (rmrjr22):

x^3 = 0 , (x-2)^2 = 0 , (x-1) = 0

OpenStudy (rmrjr22):

get it?